# Beyond Sociodemographic Prompting: Using Supervision to Align LLMs with Human Response Distributions

### v1

@@ 6.2 How are individual SD groups affected?
6.3 Do models generalize out-of-domain?
We also study whether our regression models for calibration generalize to unseen datasets, and whether calibrated distributions are more aligned than the original LLM-generated distributions for the unseen dataset. We train regression models on two of our three datasets and evaluate on the held-out dataset. Results are shown in Appendix A.5. As expected, opinion alignment is lower on unseen dataset, but the distributions calibrated on out-of-distribution data are typically more aligned with human responses than the original LLM-generations of that dataset. Alignment is higher in 92.8% of settings for the unseen OQA dataset, 90.6% of settings for WVS, and 85.5% for WGM. This suggests that regression models trained on these datasets could generalize to data, though it would be better to collect a few supervised examples in-domain to train regression models.
7 Does post-training impact alignment?
To study effects of post-training methods for RQ5, we compare differences in performance between four OLMo-2-7B models (the only completely open source model family we evaluate): the base model, base + SFT, base + SFT + DPO, and the instruct model (base + SFT + DPO + RLVR). We study differences between base vs. SD and calibrated vs. uncalibrated with the verbalized elicitation method, since it generally is best aligned.
Findings (RQ5):
Calibrated alignment is similar both before and after each phase of post-training, though pre-calibrated alignment decreases for 2/3 datasets when post-training is added.
7.1 What is the effect of calibration?
Figure 3 plots opinion alignment for all OLMo models. Interestingly, calibration appears to increase opinion alignment less for the base model, and much more so for post-trained models on the WGM and OQA datasets. OLMo models are also much more aligned with human responses on the WVS overall (both with and without calibration) compared to the other two datasets, and calibration appears to affect opinion alignment much less.
7.2 What is the effect of post-training?
Changes in alignment after various post-training methods is dataset-dependent. For WGM, alignment of the base distribution falls significantly after post-training, though alignment for the SD distributions slightly increase with RLVR. Calibration alignment stays consistent after post-training. On OQA, alignment decreases with SFT, but stays about the same when adding other post-training. This is true of base, SD, and calibrated distributions, though calibration alignment decreases the least. On WVS however, alignment increases after post-training methods are added, and alignment changes less with calibration.
We suspect the increase in performance on WVS when adding post-training indicates that fine-tuning data might align more with global populations surveyed in WVS. The drop in performance for uncalibrated distributions on WGM and OQA might indicate that OLMO’s post-training data is less aligned with some of the populations of these datasets.
8 Suggestions for practitioners
One must acknowledge risks and and exercise caution in using LLMs to approximate responses from diverse SD groups, since fallible LLMs may naturally exhibit random errors and systematic biases in their outputs. That said, the ease, speed, and cost of automated approximation may motivate some practitioners in some settings to use LLMs to approximate human responses, provided the predictive accuracy is “good enough” for their needs. Whether this is so will depend greatly on the varying accuracy needs of different real-world uses cases, as well as predictive accuracy relative to the SD groups of interest to each use case. We also note that such approximation may represent only an initial starting point, providing an initial guess of the data while awaiting for the real human response data to more slowly trickle in from participants.
In general, we suggest that practitioners: (1) evaluate distributions rather than a single response when evaluating alignment with human responses, and (2) calibrate LLM-generated distributions. As our results shown, calibration can be effective with as few as 5 supervised examples, though degradation varies different SD groups, so one must acknowledge the tradeoffs between risk, cost and quality in performing calibration with less supervision. We provide a summary of our results for the five models shown in Table 1.
8.1 Which methods and models work best?
As shown in Table 1, although the best distribution elicitation method is LLM and dataset dependent, verbalized elicitation leads to the best aligned distributions in the majority of settings and has a higher average across settings. Although this is true both pre-calibration and post-calibration, all methods are much more closely aligned post-calibration: without calibration, verbalized elicitation is the most aligned in 90% of settings, and with calibration it is the most aligned in 63% of settings. Adding sociodemographic information does not make LLM distributions more aligned with human distributions consistently; SD verbalized distributions are more aligned than their base verbalized counterparts in 67% of settings pre-calibration and 60% of settings post-calibration.
Overall, using verbalized elicitation (followed by calibration) might be best to obtain the most aligned distributions. However, we note that after calibration, the alignment of distributions produced by all the methods are fairly close (within 2% of each other). We also find that prompting with SD information is not as important, though might lead to slightly more aligned distributions in aggregate.
The most aligned models are much more dataset dependent. Claude 3.5 v2 is most aligned for WGM and OQA pre-calibration (with and without SD information). Post-calibration, Claude is most aligned for OQA, while Llama 3.2 90B is most aligned for WGM. For WVS, OLMo 7B Instruct is most aligned (both pre/post calibration).
@@ 9 Discussion
Essentializing sociodemographics. We note that by prompting LLMs with SD information – and evaluating their generated distributions against human participants with those sociodemographics– we are essentializing the sociodemographic to the identities of the survey respondents and their opinions Wang et al. (2024). However, practitioners still seek to understand opinions of specific populations (e.g., for election polling) where an aligned LLM might be beneficial for faster iteration on survey questions. Although sociodemographics are not at all the only factor that influences people’s opinions, they certainly play a part Sap et al. (2022); Biester et al. (2022); Pei and Jurgens (2023).
@@
The effect of SD prompting. While our SD information provides valuable context about us, it alone does not determine our opinions, and it is fallacious to assume otherwise. As we have seen, prompting with SD information on its own also does not necessarily lead to more aligned distributions. Future work might study eliciting distributions with more implicit demographic information or with past opinions as prior work has done with single responses Hwang et al. (2023); Do et al. (2025). One could make personalized predictions for a sample of a group’s individuals members Gordon et al. (2022) rather than a single distributive prediction for the group. Multimodal prompts might also add SD information implicitly, though this might come with additional biases.
@@
LLM distributions are uncalibrated. We find that LLM-generated distributions (produced in answer to survey questions) are uncalibrated – simple regression allows us to scale distributions to be much more aligned with human opinion in aggregate, though this necessarily is more aligned with some populations over others. We also find that calibration reduces variance, both across examples and across settings. Since calibrated distributions generally have higher performance, this might provide evidence for LLMs exaggerating differences between sociodemographic groups; exaggerations that are reduced with calibration.
@@
Calibration on individual SD groups. We do not find many significant differences between calibrated/uncalibrated distributions at the individual SD group level. Calibrating responses from each individual SD would likely improve alignment for that particular SD, which we do not do here. However, we caution efficacy may vary by demographic; some demographic traits might be more relevant to the response distributions of certain questions (e.g., urban vs. rural areas might affect gun rights questions more than marital status).
@@
The dataset, LLM, and distribution elicitation method matter. As prior work has found when evaluating SD prompting on human majority responses Beck et al. (2024); Hu and Collier (2024), in this work we find that opinion alignment is dataset, LLM, and distribution elicitation method dependent. However, we find that verbalized elicitation tends to elicit probabilities that are most aligned with human distributions in the most settings, both before and after calibration. We find that the Claude 3.5 models are most aligned with human responses on average across all datasets, though we highlight that this is dataset-dependent.
@@ 10 Conclusion
We investigate LLM distribution alignment with human responses to subjective large-scale surveys, both on average and across diverse population groups. We show that using simple supervision can improve alignment with population groups consistently across datasets, models, and distribution elicitation techniques. Our work also offers guidance for those using LLMs to predict human responses in practice, and our benchmark can enable and stimulate future research.
@@ 11 Limitations
Prompting methods. We do not study chain of thought or other prompting methods which might increase overall performance. We do not experiment with varied temperatures for our generations. For elicitation methods, though a larger might improve generated distributions, our choice of is reasonable given inference costs. We use for verbalized elicitation (as opposed to ) since we found little variance between runs. We also leave in-context learning or fine-tuning Orlikowski et al. (2025) with human responses to future work.
Logit-based elicitation methods. We focus on methods that work most broadly, supporting use with both black-box and white-box LLMs. That said, we do also report a smaller scale study with logit-based distribution elicitation with four Llama models (Appendix A.3). We find that using log probabilities leads to the most aligned distributions in 10/24 or 42% of settings with calibration. Future work might look more into comparisons between logit-based and verbalized distribution elicitation.
Demographics studied. We only studied a subset of the demographics available in the survey datasets, and each demographic individually. Future work might study a larger subset of demographics, as well as intersectional demographics.
Regression. Our regression predicts each answer choice individually and then normalizes the predicted answers. We also optimize regression for MSE instead of the alignment metric. We tried constraining optimization to learn weights for all answer choices simultaneously and enforce proper distributions, but learning the answer choices individually performed better. Future work might also ensemble multiple elicitation methods.
Ethics Statement
In this work, we study how we might align LLMs to the survey response distributions of various social groups. While this can be valuable for social good – for practitioners, NLP researchers, or social science researchers seeking to accurately approximate human responses for survey development, annotation tasks, or social science studies – this can also be used for adversarial purposes, such as chat bots attempting to simulate different user groups, targeted advertising, or targeted misinformation.
We also note that evaluating LLM responses against human responses of particular sociodemographic groups often assumes that everyone in a particular demographic group thinks the same way, and essentializes the demographic to an individual’s identity. In this work, we try to address this by evaluating against the distribution of human responses for a particular sociodemographic group. By doing so, we hope to better align LLMs with various sociodemographic groups in a distributionally pluralistic manner, rather than assume that all members of the group would answer the same way.
@@ Acknowledgments
We thank Sooyong Lee for early contributions that did not make it into the ultimate paper. This research was supported in part by Cisco, Good Systems222https://goodsystems.utexas.edu/ (a UT Austin Grand Challenge dedicated to developing responsible AI technologies), and a grant from Open Philanthropy. The statements made herein are solely the opinions of the authors and do not reflect the views of the sponsoring agencies.
@@ A.1 Alignment results for all models
A.2 Minimal Supervision
Results are in Figure 4.
A.3 Log Probability Results
We get the model’s log probabilities for each answer choice Geng et al. (2024b) and normalize to get the distribution over all answer choices, following prior work Santurkar et al. (2023). We obtain log probabilities for the smaller Llama models. Log probability results are shown in Table 4.
@@ A.3 Log Probability Results
A.4 Individual Sociodemographic Results
Results for all demographics for the WGM, OQA, and WVS datasets are in Tables 6, 7, and 8 respectively.
@@ A.4 Individual Sociodemographic Results
A.5 Generalization results
@@ B.1 Demographics used
| Dataset | Example Questions | Demographics |
| WGM | How much do you, personally, know about science? Do you know a lot, some, not much, or nothing at all? | age, education, employment status, gender, income quintile, living area (city or rural), world region |
| How much do you trust doctors and nurses in your country? | ||
| OQA | How important, if at all, is being a gun owner to your overall identity? | age, education, income, marital status, political ideology, political party, race, US region, religion, sex |
| How much confidence, if any, do you have in elected officials to act in the best interests of the public? | ||
| WVS | Do you agree strongly, agree, disagree, or disagree strongly with the following statement? One of my main goals in life has been to make my parents proud. | age, education, employment status, household size, immigrant status, income, marital status, religion, sex |

### v2

@@ 8 Discussion
Limitations of demographic modeling. Modeling groups wrt. their demographics assuredly provides only a narrow, incomplete view of people, and demographics alone do not determine our opinions. That said, such models may usefully approximate distributions of group opinions, provided care is exercised. We note that our methods for improving alignment should not be viewed as a way to replace human respondents in surveys, but rather a step towards more accurately representing groups.
@@
The effect of SD prompting. We showed that simply prompting with SD information does not necessarily lead to more aligned distributions. Future work might study eliciting distributions with more implicit demographic information or with past opinions as prior work has done with single responses Hwang et al. (2023); Do et al. (2025). One could make personalized predictions for a sample of a group’s individuals members Gordon et al. (2022); Hwang et al. (2023) rather than a single distributive prediction for the group. Multimodal prompts might also add SD information implicitly, though this might come with additional biases.
@@
LLM distributions are uncalibrated. We find that LLM-generated distributions (produced in answer to survey questions) are uncalibrated; simple regression allows us to scale distributions to be more aligned with human opinion in aggregate, though this necessarily is more aligned with some populations over others. We also find that calibration reduces variance, both across examples and across settings. Since calibrated distributions generally have higher performance, this suggests LLMs exaggerate differences between sociodemographic groups, which calibration might help mitigate.
@@
Calibration on individual SD groups. We do not find many significant differences between calibrated/uncalibrated distributions at the individual SD group level. However, we hypothesize that calibrating responses from each individual SD would likely improve its alignment, though we leave this to future work. However, efficacy may vary by demographic; some demographic traits are more relevant to the response distributions of certain questions (e.g., urban vs. rural areas might affect gun rights questions more than marital status).
@@
The choice of LLM and distribution elicitation method matters less with calibration. We find that distributional opinion alignment is dataset, LLM, and distribution elicitation method dependent, consistent with prior work evaluating majority responses Beck et al. (2024). However, verbally-elicited distributions are most aligned with human distributions in the most settings. We also find that the Claude 3.5 models are most aligned with human responses on average across all datasets, though we highlight that this is also dataset-dependent. More importantly however, we find that calibration reduces variance in alignment across models and methods; i.e., the choice of model and elicitation method matters less after calibration.
@@ 9 Conclusion
We investigate LLM distribution alignment with human responses to subjective large-scale surveys, both on average and across diverse population groups. We show that using simple supervision can improve alignment with population groups consistently across datasets, models, and distribution elicitation techniques, as well as increase consistency between settings. Our benchmark also serves to help enable and stimulate future research.
@@ 10 Limitations
Prompting methods. We do not experiment with varied temperatures for our generations. For elicitation methods, though a larger might improve generated distributions, our choice of is reasonable given inference costs and following prior work Xiong et al. (2024). We use for verbalized elicitation (as opposed to ) since we found little variance between runs.
Logit-based elicitation methods. We focus on methods that broadly work with open-source, open-model, and API-access LLMs. That said, we report a smaller scale study with logit-based distribution elicitation with four Llama models (Appendix A.2). We found that log probabilities with calibration yielded the most aligned distributions in 10/24 settings considered. Future work might look more into comparisons between logit-based and verbalized distribution elicitation.
Demographics studied. We only studied a subset of the demographics available in the datasets. Future work might study additional and/or intersectional demographics Crenshaw (1991).
Regression. Our regression predicts each answer choice individually and then normalizes the predicted answers. We tried constraining optimization to learn weights for all answer choices simultaneously and enforce proper distributions, but learning the answer choices individually performed better. We also optimize regression for MSE instead of the alignment metric. Future work might also ensemble multiple elicitation methods.
Ethical Considerations
In this work, we study how we might align LLMs to the survey response distributions of various social groups. While this can be valuable for creating LLMs that more accurately portray and reason about population groups, this can also be used for adversarial purposes, such as chat bots attempting to simulate different user groups, targeted advertising, or targeted misinformation.
We note that by prompting LLMs with SD information, and evaluating their distributions against humans with those SDs, we are essentializing the SD to the identities of the survey respondents and their opinions Wang et al. (2025a). Although SDs are not at all the only factor that influences people’s opinions, they certainly play a part Sap et al. (2022); Biester et al. (2022); Pei and Jurgens (2023). By seeking to infer distributions of human responses rather than majority responses for each SD group, we seek to align LLMs with groups in a distributionally pluralistic manner, rather than assume all members of the group would answer similarly.
@@ Acknowledgments
We thank Sooyong Lee for early contributions that did not make it into the ultimate paper. This research was supported in part by Cisco, Good Systems222https://goodsystems.utexas.edu/ (a UT Austin Grand Challenge dedicated to developing responsible AI technologies), and a grant from Open Philanthropy. We thank the Texas Advanced Computing Center (TACC) for use of its superb infrastructure. The statements made herein are solely the opinions of the authors and do not reflect the views of the sponsoring agencies.
@@ A.1 Alignment results for all models
A.2 Log Probability Results
We get the model’s log probabilities for each answer choice Geng et al. (2024) and normalize to get the distribution over all answer choices, following prior work Santurkar et al. (2023). We obtain log probabilities for the smaller Llama models. Log probability results are shown in Table 6.
@@ A.2 Log Probability Results
A.3 Individual Sociodemographic Results
Results for all demographics for the WGM, OQA, and WVS datasets are in Tables 8, 9, and 10 respectively.
@@ A.3 Individual Sociodemographic Results
A.4 Out-of-domain generalization
We study whether our regression models for calibration generalize to unseen datasets, and whether calibrated distributions are more aligned than the original LLM-generated distributions for the unseen dataset. We train regression models on two of our three datasets and evaluate on the held-out dataset. Results are shown in Tables 11 and 12. As expected, opinion alignment is lower on unseen dataset, but the distributions calibrated on out-of-distribution data are typically more aligned with human responses than the original LLM-generations of that dataset. Alignment is higher in 92.8% of settings for the unseen OQA dataset, 90.6% of settings for WVS, and 85.5% for WGM. This suggests that regression models trained on these datasets could generalize to data, though it would be better to collect a few supervised examples in-domain to train regression models.
@@ B.1 Demographics used
| Dataset | Qs | Example question | Demographics |
| WGM | 14 |
How much do you, personally, know about science? Do you know a lot, some, not much, or nothing at all?
How much do you trust doctors and nurses in your country? |
age (15-29, 30-49, 50+), education(primary, secondary, tertiary), employment status (employed full time for an employer, employed full time for self, employed part time but do not want full time, unemployed, employed part time want full time, out of work force), income quintile (poorest 20%, second 20%, middle 20%, fourth 20%, top 20%), living area (city, rural), sex (male, female), world region (Eastern Africa, Central Africa, North Africa, Southern Africa, Western Africa, Central America and Mexico, Northern America, South America, Central Asia, East Asia, Southeast Asia, South Asia, Middle East, Eastern Europe, Northern Europe, Southern Europe, Western Europe, Aus/NZ) |
| OQA | 38 |
How important, if at all, is being a gun owner to your overall identity?
How much confidence, if any, do you have in elected officials to act in the best interests of the public? |
age (18-29, 30-49, 50-64, 65+), education (Less than high school, High school graduate, Some college, no degree, Associates degree, College graduate/some postgrad, Postgraduate), income (Less than $30,000, $30,000-$50,000, $50,000-$75,000, $75,000-$100,000, $100,000 or more), marital status (Married, Divorced, Separated, Widowed, Never been married), political ideology (Very conservative, Conservative, Moderate, Liberal, Very liberal), political party (Republican, Democrat, Independent, Other), race (White, Black, Asian, Hispanic, Other), US region (Northeast, Midwest, South, West), religion (Protestant, Roman Catholic, Mormon, Orthodox, Jewish, Muslim, Buddhist, Hindu, Atheist, Agnostic, Other, Nothing in particular), sex (Male, Female) |
| WVS | 40 |
Do you agree strongly, agree, disagree, or disagree strongly with the following statement? One of my main goals in life has been to make my parents proud.
@@
age (16-24 years, 25-34 years, 35-44 years, 45-54 years, 55-64 years, 65+ years), education (early childhood, primary, lower secondary, upper secondary, post-secondary, short-cycle tertiary, bachelor or equivalent, master or equivalent, doctoral or equivalent), employment status (full time, part time, self-employed, retired/pensioned, housewife not otherwise employed, student, unemployed), household size (1, 2, 3, 4, 5, 6, 7 persons or more), immigrant status (I am born in this country, I am an immigrant to this country), income (low, middle, high), marital status (Married, Living together as married, Divorced, Separated, Widowed, Single), religion (none, Roman Catholic, Protestant, Orthodox (Russian/Greek/etc.), Jew, Muslim, Hindu, Buddhist, Other Christian), sex (Male, Female) |
@@ B.3 Sociodemographic prompts
B.4 Prompt variations
Prompt variations are shown in Table 16.
| Prompt variation | Prompt template |
| Few shot | Please complete the task below. Follow the examples given. Examples: {5 examples} Task: {Original prompt} |
| CoT | {Original question} Explain your reasoning step-by-step before answering. {Response instructions} Put your final answer at the end of your response between <ANS_START> and </ANS_END>. Use a maximum of 500 words for your reasoning and final answer combined. |
Appendix C Extracted distributions
The percent of distributions extracted from each LLM for each dataset and elicitation method is shown in Table 17.
| Model | Base prompt | SD prompt | |||||
| WGM | Claude-3 | 100 | 100 | 100 | 100 | 100 | 100 |
| Claude-3.5-v1 | 100 | 100 | 100 | 100 | 100 | 100 | |
| Claude-3.5-v2 | 100 | 100 | 100 | 100 | 100 | 100 | |
| Llama-3-70B | 100 | 100 | 100 | 100 | 100 | 100 | |
| Llama-3.1-70B | 100 | 100 | 82.4 | 100 | 100 | 92.8 | |
| Llama-3.2-1B | 100 | 100 | 0 | 100 | 100 | 1.3 | |
| Llama-3.2-11B | 100 | 100 | 70.6 | 100 | 100 | 27.1 | |
| Llama-3.2-90B | 100 | 100 | 100 | 100 | 100 | 94.9 | |
| Mistral-small | 100 | 100 | 100 | 100 | 100 | 99.7 | |
| Mistral-large | 100 | 100 | 100 | 100 | 100 | 100 | |
| OLMo-2-7B-B | 100 | 100 | 47.1 | 100 | 100 | 69.3 | |
| OLMo-2-7B-S | 100 | 100 | 100 | 100 | 100 | 97.6 | |
| OLMo-2-7B-D | 100 | 100 | 100 | 100 | 100 | 92.9 | |
| OLMo-2-7B-I | 100 | 100 | 70.6 | 100 | 100 | 89.1 | |
| Qwen-2.5-72B | 100 | 100 | 100 | 100 | 100 | 100 | |
| OQA | Claude-3 | 100 | 100 | 100 | 100 | 100 | 100 |
| Claude-3.5-v1 | 100 | 100 | 100 | 100 | 100 | 100 | |
| Claude-3.5-v2 | 100 | 100 | 100 | 100 | 100 | 100 | |
| Llama-3-70B | 100 | 100 | 91.9 | 100 | 100 | 99.8 | |
| Llama-3.1-70B | 100 | 100 | 100 | 100 | 12.1 | 93.4 | |
| Llama-3.2-1B | 100 | 100 | 0 | 100 | 100 | 0.6 | |
| Llama-3.2-11B | 100 | 100 | 32.4 | 100 | 100 | 7.2 | |
| Llama-3.2-90B | 100 | 100 | 100 | 100 | 100 | 99.7 | |
| Mistral-small | 100 | 100 | 100 | 100 | 100 | 100 | |
| Mistral-large | 100 | 100 | 100 | 100 | 100 | 99.7 | |
| OLMo-2-7B-B | 100 | 100 | 100 | 100 | 100 | 99.9 | |
| OLMo-2-7B-S | 100 | 100 | 100 | 100 | 100 | 99.5 | |
| OLMo-2-7B-D | 100 | 100 | 94.6 | 100 | 100 | 79.8 | |
| OLMo-2-7B-I | 100 | 100 | 75.7 | 100 | 100 | 79.5 | |
| Qwen-2.5-72B | 100 | 100 | 100 | 100 | 100 | 99.4 | |
| WVS | Claude-3 | 100 | 100 | 100 | 100 | 100 | 100 |
| Claude-3.5-v1 | 100 | 100 | 100 | 19.6 | 19.6 | 19.6 | |
| Claude-3.5-v2 | 100 | 100 | 100 | 49 | 49 | 49 | |
| Llama-3-70B | 100 | 100 | 97.9 | 99.4 | 99 | 87.2 | |
| Llama-3.1-70B | 100 | 100 | 100 | 49 | 49 | 47.2 | |
| Llama-3.2-1B | 100 | 100 | 0 | 49 | 49 | 0.4 | |
| Llama-3.2-11B | 100 | 100 | 39.7 | 49 | 49 | 3.7 | |
| Llama-3.2-90B | 100 | 100 | 100 | 100 | 100 | 98.3 | |
| Mistral-small | 100 | 100 | 100 | 100 | 100 | 100 | |
| Mistral-large | 100 | 100 | 100 | 100 | 100 | 99.9 | |
| OLMo-2-7B-B | 100 | 100 | 100 | 100 | 100 | 99.6 | |
| OLMo-2-7B-S | 100 | 100 | 100 | 100 | 100 | 99.3 | |
| OLMo-2-7B-D | 100 | 100 | 76.8 | 100 | 100 | 72.2 | |
| OLMo-2-7B-I | 100 | 100 | 82.1 | 100 | 100 | 79.3 | |
| Qwen-2.5-72B | 100 | 100 | 100 | 100 | 100 | 98.6 | |